# Lindblad dynamics from spatio-temporal correlation functions in nonintegrable spin-$1/2$ chains with different boundary conditions

### v1

@@ I Introduction
In an open-system scenario, transport can be induced by the coupling of the system to baths at different temperatures or chemical potentials, which then usually yields a nonequilibrium steady state (NESS) with a constant current and characteristic density profile in the long-time limit [24, 25, 26, 27]. Such a scenario is often modelled by a equation of Lindblad form [28]. While the derivation of this equation from a microscopic system-bath model is a nontrivial task in practice [26, 29], it is the most general version of a time-local master equation, which maps any density matrix to a density matrix again. In particular, it allows for an efficient numerical treatment by matrix-product states for quite large system sizes [24, 30, 31, 32], as dissipation reduces the unavoidable growth of entanglement as a function of time.
@@ II Closed models and spatio-temporal correlation functions
$H_{\text{pbc}}=H_{\text{obc}}+J(S_{N}^{x}S_{1}^{x}+S_{N}^{y}S_{1}^{y}+\Delta S_{N}^{z}S_{1}^{z})\,.$ (2)
The specific choice of boundary conditions will play an important role for the open setup, which will be discussed later in detail.
@@
Figure 2: (a) Time-space density plot of the correlation function $\langle S_{r}^{z}(t)S_{r^{\prime}}^{z}(0)\rangle_{\text{eq}}$, as obtained numerically for the Hamiltonian $H^{\prime}$ in Eq. (3) with $\Delta=1.5$, $\Delta^{\prime}=0.5$, $N=20$, and periodic boundary conditions. (b) Corresponding time evolution of the standard deviation $\Sigma(t)$ as well as two power laws $\propto t^{1}$ and $\propto t^{1/2}$, in a double-logarithmic plot.
@@ V Stochastic unraveling of the Lindblad equation
Therefore, the dynamics of a pure state is only generated by the closed system $H$, apart form the scalar damping term. This fact fact will be one of the main ingredients to connect the closed system and the weakly driven open system. For larger $\mu$, the dynamics is more complicated and also involves the operators $n_{B_{1}}$ and $n_{B_{2}}$.
@@
This sequence of stochastic jumps and deterministic evolutions leads to a particular trajectory $|\psi_{T}(t)\rangle$. The time-dependent density matrix according to the Lindblad equation can be eventually approximated by the average over different trajectories $T$. Thus, the expectation value read
@@ VI.1 Time evolution
Figure 3: Temporal decay of the equal-site correlation function $\langle S_{r}^{z}(t)S_{r}^{z}(0)\rangle_{\text{eq}}$ for (a) periodic boundary conditions and (b) open boundary conditions, as obtained numerically for the Hamiltonian $H^{\prime}$ in Eq. (3) for different $\Delta$ at fixed $\Delta^{\prime}=0.5$ and $N=20$. The equilibration value $1/(4N)$ and a time scale $1/(2\gamma)$ with $\gamma/J=0.1$ are indicated.
@@
$\tau_{j+1}=\tau_{j}-\frac{\varepsilon_{j+1}}{2\gamma}\,,$ (36)
@@
In principle, the prediction in Eqs. (32) and (35) can be calculated analytically for a specific model. However, the closed-system correlation functions $\langle S_{r}^{z}(t)S_{B_{1}}^{z}(0)\rangle_{\text{eq}}$ and $\langle S_{r}^{z}(t)S_{B_{2}}^{z}(0)\rangle_{\text{eq}}$ are often available only numerically, such that the prediction has to calculated numerically as well.
@@ VI.2 Injected magnetization
While Eqs. (32) and (35) allow to predict the dynamics of magnetization at finite times and in the limit of long times, a similar expression can derived for the respective currents in the steady state. To this end, let us consider the magnetization injected by the first bath, can be predicted as [52]
@@
$\langle j_{B_{1}}\rangle=\frac{\text{d}}{\text{d}t}\frac{\langle S_{B_{1}}^{z}(t)\rangle}{f}\,.$ (40)
@@ VII.2 Open boundary conditions and different bath-coupling geometries
Figure 9: On a bath coupling close to the edges. (a) Temporal decay of the equal-site correlation function $\langle S_{r}^{z}(t)S_{r}^{z}(0)\rangle_{\text{eq}}$ for different sites $r$, as obtained numerically for the model $H^{\prime}$ in Eq. (3) with $\Delta=1.5$, $\Delta^{\prime}=0.5$, $N=20$, and open boundary conditions. (b) Open-system dynamics for a bath coupling at sites $B_{1}=3$ and $B_{2}=17$. In comparison to the data in Fig. 7 (c), the agreement between the prediction (Eq.) and stochastic unraveling (data) is better.
@@
For $\Delta=0.5$, where the assumption is fulfilled best, we increase next the system size to $N=34\gg 20$. For such a system size, the Hilbert-space dimension becomes huge and stochastic unraveling is not feasible any more, due to the many trajectories required. However, the prediction can still be carried out, since the two correlation functions $\langle S_{r}^{z}(t)S_{B_{1}}^{z}(0)\rangle_{\text{eq}}$ and $\langle S_{r}^{z}(t)S_{B_{2}}^{z}(0)\rangle_{\text{eq}}$ need to be calculated only once. In particular, this calculation is possible by the use of dynamical quantum typicality and supercomputers. In Fig. 8, we depict the corresponding prediction and compare to the TEBD solution, which up to times $tJ\approx 50$ does not depend on the bond dimension used. (A convergence analysis of TEBD can be found in the appendix.) The convincing agreement supports that our prediction is an useful alternative for large system sizes, which are usually only accessible by matrix product states.
@@ VII.3 Staggered field
Figure 10: Another model: staggered field. (a) Temporal decay of the equal-site correlation function $\langle S_{r}^{z}(t)S_{r}^{z}(0)\rangle_{\text{eq}}$ for closed and open boundary conditions, as obtained numerically for the Hamiltonian $H^{\prime\prime}$ in Eq. (4) with $\Delta=1.0$, $B/J=0.5$, and $N=20$. (b) and (c) Open-system dynamics for the respective boundary conditions as well as $\gamma/J=0.1$ and $\mu=0.1$. The prediction (Eq.) is compared to either stochastic unraveling (data) or time-evolving block decimation (TEBD).
@@ VIII Conclusion
To summarize, we have studied the Lindblad equation as a central approach to boundary-driven magnetization transport in spin-$1/2$ chains. Our main motivation has been to understand to what extend the dynamics in the open system, at finite times and in the limit of long times, can be predicted on the basis of the dynamics in the closed system. To this end, we have followed the idea of a previous work [52], which has suggested a prediction in terms of spatio-temporal correlation functions, Eqs. (32) and (35), given the case of weak driving and small system-bath coupling. While this work was focused on integrable systems and periodic boundary conditions, we have substantially extended the analysis in the current work by going in three different directions: We have (i) considered nonintegrable systems, (ii) included open boundary conditions and other bath-coupling geometries, and (iii) provided a comparison with time-evolving block decimation.
@@
Promising future directions of future research include quasi-1D lattices, finite temperatures, energy transport, fermionic models, and disorder.
Figure 12: (a) Dynamics of the equal-site correlation function $\langle S_{r}^{z}(t)S_{r}^{z}(0)\rangle_{\text{eq}}$ for open boundary conditions, as depicted in Fig. 3 (b) for different $\Delta$, but now for two system sizes $N=20$ and $N=24$. (b) Time-space density plot of the correlation functions $\langle S_{r}^{z}(t)S_{r^{\prime}}^{z}(0)\rangle_{\text{eq}}$ for $N=34$ and $\Delta=0.5$, which are used for the prediction in Fig. 8.
@@ Acknowledgments
Our work has been funded by the Deutsche Forschungsgemeinschaft (DFG), projects 397107022 (GE 1657/3-2), 397300368 (MI 1772/4-2), and 397067869 (STE 2243/3-2), within DFG Research Unit FOR 2692, grant no. 355031190. J. R. acknowledges funding from the European Union’s Horizon Europe research and innovation programme, Marie Sklodowska-Curie grant no. 101060162, and the Packard Foundation through a Packard Fellowship in Science and Engineering. We gratefully acknowledge the Gauss Centre for Supercomputing e. V. for funding this project by providing computing time on the GCS Supercomputer JUWELS6 at Jülich Supercomputing Centre (JSC). S. N. acknowledges support by the projects J1-2463 and P1-0044 program of the Slovenian Research Agency, EU via QuantERA grant T-NiSQ, and also computing time for the TEBD calculations at the supercomputer Vega at the Institute of Information Science (IZUM) in Maribor, Slovenia.
@@ Appendix B Correlation functions for open boundary conditions
For completeness, Fig. 12 (b) shows the full time-space dependence of the correlation functions $\langle S_{r}^{z}(t)S_{r^{\prime}}^{z}(0)\rangle_{\text{eq}}$ for $N=24$ and $\Delta=0.5$, which have been used for the prediction in Fig. 8.

### v2

@@ I Introduction
In an open-system scenario, transport can be induced by the coupling of the system to baths at different temperatures or chemical potentials, which then usually yields a nonequilibrium steady state (NESS) with a constant current and characteristic density profile in the long-time limit [24, 25, 26, 27]. Such a scenario is often modelled by an equation of Lindblad form [28]. While the derivation of this equation from a microscopic system-bath model is a nontrivial task in practice [26, 29], it is the most general version of a time-local master equation, which maps any density matrix to a density matrix again. In particular, it allows for an efficient numerical treatment by matrix-product states for quite large system sizes [24, 30, 31, 32], as dissipation reduces the unavoidable growth of entanglement as a function of time.
@@ II Closed models and spatio-temporal correlation functions
$H_{\text{pbc}}=H_{\text{obc}}+J(S_{N}^{x}S_{1}^{x}+S_{N}^{y}S_{1}^{y}+\Delta S_{N}^{z}S_{1}^{z})\,,$ (2)
where the numbers of sites $N$ is chosen to be even. The specific choice of boundary conditions will play an important role for the open setup, which will be discussed later in detail.
@@
Figure 2: (a) Time-space density plot of the correlation function $\langle S_{r}^{z}(t)S_{r^{\prime}}^{z}(0)\rangle_{\text{eq}}$, as obtained numerically for the Hamiltonian $H^{\prime}$ in Eq. (3) with $\Delta=1.5$, $\Delta^{\prime}=0.5$, $N=20$, $r^{\prime}=N/2+1=11$, and periodic boundary conditions. (b) Corresponding time evolution of the standard deviation $\Sigma(t)$ as well as two power laws $\propto t^{1}$ and $\propto t^{1/2}$, in a double-logarithmic plot.
@@ V Stochastic unraveling of the Lindblad equation
Therefore, the dynamics of a pure state is only generated by the closed system $H$, apart form the scalar damping term. This fact will be one of the main ingredients to connect the closed system and the weakly driven open system. For larger $\mu$, the dynamics is more complicated and also involves the operators $n_{B_{1}}$ and $n_{B_{2}}$.
@@
This sequence of stochastic jumps and deterministic evolutions leads to a particular trajectory $|\psi_{T}(t)\rangle$. The time-dependent density matrix according to the Lindblad equation can be eventually approximated by the average over different trajectories $T$. Thus, the expectation value reads
@@ VI.1 Time evolution
Figure 3: Temporal decay of the autocorrelation function $\langle S_{1}^{z}(t)S_{1}^{z}(0)\rangle_{\text{eq}}$ for (a) periodic boundary conditions and (b) open boundary conditions, as obtained numerically for the Hamiltonian $H^{\prime}$ in Eq. (3) for different $\Delta$ at fixed $\Delta^{\prime}=0.5$ and $N=20$. The equilibration value $1/(4N)$ and a time scale $1/(2\gamma)$ with $\gamma/J=0.1$ are indicated.
@@
$\tau_{j+1}=\tau_{j}-\ln\frac{\varepsilon_{j+1}}{2\gamma}\,,$ (36)
@@
In principle, the prediction in Eqs. (32) and (35) can be calculated analytically for a specific model. However, the closed-system correlation functions $\langle S_{r}^{z}(t)S_{B_{1}}^{z}(0)\rangle_{\text{eq}}$ and $\langle S_{r}^{z}(t)S_{B_{2}}^{z}(0)\rangle_{\text{eq}}$ are often available only numerically, such that the prediction has to be calculated numerically as well.
@@ VI.2 Injected magnetization
While Eqs. (32) and (35) allow to predict the dynamics of magnetization at finite times and in the limit of long times, a similar expression can derived for the respective currents in the steady state. To this end, let us consider the magnetization injected by the first bath, which can be predicted as [52]
@@
$\langle j_{B_{1}}\rangle=\frac{\text{d}}{\text{d}t}\frac{\langle\delta S_{B_{1}}^{z}(t)\rangle}{f}\,.$ (40)
@@ VII.2 Open boundary conditions and different bath-coupling geometries
Figure 9: On a bath coupling close to the edges. (a) Temporal decay of the autocorrelation function $\langle S_{r}^{z}(t)S_{r}^{z}(0)\rangle_{\text{eq}}$ for different sites $r$, as obtained numerically for the model $H^{\prime}$ in Eq. (3) with $\Delta=1.5$, $\Delta^{\prime}=0.5$, $N=20$, and open boundary conditions. (b) Open-system dynamics for a bath coupling at sites $B_{1}=3$ and $B_{2}=17$. In comparison to the data in Fig. 7 (c), the agreement between the prediction (Eq.) and stochastic unraveling (data) is better.
@@
For $\Delta=0.5$, where the assumption is fulfilled best, we increase next the system size to $N=34\gg 20$. For such a system size, the Hilbert-space dimension becomes huge and stochastic unraveling is not feasible any more, due to the many trajectories required. However, the prediction can still be carried out, since the two correlation functions $\langle S_{r}^{z}(t)S_{B_{1}}^{z}(0)\rangle_{\text{eq}}$ and $\langle S_{r}^{z}(t)S_{B_{2}}^{z}(0)\rangle_{\text{eq}}$ need to be calculated only once. In particular, this calculation is possible by the use of dynamical quantum typicality and supercomputers. In Fig. 8, we depict the corresponding prediction and compare to the TEBD solution, which up to times $tJ\approx 50$ does not depend on the bond dimension used. (A convergence analysis of TEBD can be found in the appendix.) The convincing agreement supports that our prediction is a useful alternative for large system sizes, which are usually only accessible by matrix product states.
@@ VII.3 Staggered field
Figure 10: Another model: staggered field. (a) Temporal decay of the autocorrelation function $\langle S_{1}^{z}(t)S_{1}^{z}(0)\rangle_{\text{eq}}$ for closed and open boundary conditions, as obtained numerically for the Hamiltonian $H^{\prime\prime}$ in Eq. (4) with $\Delta=1.0$, $B/J=0.5$, and $N=20$. (b) and (c) Open-system dynamics for the respective boundary conditions as well as $\gamma/J=0.1$ and $\mu=0.1$. The prediction (Eq.) is compared to either stochastic unraveling (data) or time-evolving block decimation (TEBD).
@@ VIII Conclusion
To summarize, we have studied the Lindblad equation as a central approach to boundary-driven magnetization transport in spin-$1/2$ chains. Our main motivation has been to understand to what extent the dynamics in the open system, at finite times and in the limit of long times, can be predicted on the basis of the dynamics in the closed system. To this end, we have followed the idea of a previous work [52], which has suggested a prediction in terms of spatio-temporal correlation functions, Eqs. (32) and (35), given the case of weak driving and small system-bath coupling. While this work was focused on integrable systems and periodic boundary conditions, we have substantially extended the analysis in the current work by going in three different directions: We have (i) considered nonintegrable systems, (ii) included open boundary conditions and other bath-coupling geometries, and (iii) provided a comparison with time-evolving block decimation.
@@
Promising future directions of future research include quasi-1D lattices, finite temperatures, energy transport, fermionic models, and disorder. Another interesting avenue is to explore in which cases the slowly decaying edge modes of the closed system can be enhanced in the open system [65].
Figure 12: (a) Dynamics of the autocorrelation function $\langle S_{1}^{z}(t)S_{1}^{z}(0)\rangle_{\text{eq}}$ for open boundary conditions, as depicted in Fig. 3 (b) for different $\Delta$, but now for two system sizes $N=20$ and $N=24$. (b) Time-space density plot of the correlation functions $\langle S_{r}^{z}(t)S_{1}^{z}(0)\rangle_{\text{eq}}$ for $N=34$ and $\Delta=0.5$, which are used for the prediction in Fig. 8.
@@ Acknowledgments
Our work has been funded by the Deutsche Forschungsgemeinschaft (DFG), projects 397107022 (GE 1657/3-2), 397300368 (MI 1772/4-2), and 397067869 (STE 2243/3-2), within DFG Research Unit FOR 2692, grant no. 355031190. J. R. acknowledges funding from the European Union’s Horizon Europe research and innovation programme, Marie Sklodowska-Curie grant no. 101060162, and the Packard Foundation through a Packard Fellowship in Science and Engineering. We gratefully acknowledge the Gauss Centre for Supercomputing e. V. for funding this project by providing computing time on the GCS Supercomputer JUWELS6 at Jülich Supercomputing Centre (JSC). S. N. was supported by QuantERA grants QuSiED and T-NiSQ, by MVZI, QuantERA II JTC 2021. TEBD calculations were performed on the supercomputer Vega at the Institute of Infor- mation Science (IZUM) in Maribor, Slovenia.
@@ Appendix B Correlation functions for open boundary conditions
For completeness, Fig. 12 (b) shows the full time-space dependence of the correlation functions $\langle S_{r}^{z}(t)S_{r^{\prime}}^{z}(0)\rangle_{\text{eq}}$ for $N=34$ and $\Delta=0.5$, which have been used for the prediction in Fig. 8.